OpenStudy (anonymous):

Solve the system by substitution. {-x-y-z=-8 {-4x+4y+5z=7 {2x+2z=4

jigglypuff314 (jigglypuff314):

first isolate x in the third equation so 2x + 2z = 4 => 2x = -2z + 4 => x= -z + 2 then you can plug x into the first equation and reduce so -x -y -z = -8 => -(-z+2)-y-z=-8 => z-2-y-z=-8 => y=6 substitute x and y into the second equation so -4x+4y+5z=7 => -4(-z+2)+4(6)+5z=7 => 4z-8+24+5z=7 => 9z=-9 => z=-1 then you can substitute z into the third equation to find x 2x+2z=4 => 2x+2(-1)=4 => 2x-2=4 => 2x=6 => x=3 so x=3, y=6, and z=-1

OpenStudy (anonymous):

Thanks!

OpenStudy (wolf1728):

Just to make things easier, I multiplied the first equation by -1 A) x + y + z = 8 B) -4x + 4y + 5z = 7 C) 2x + 2z = 4 equation C can be changed to x = 2 -z Substituting this into equation A 2 - z + y + z = 8 equals y = 6 we can substitute this into equation 1 and 2 A) x + 6 + z = 8 B) -4x + 24 + 5z = 7 OR B) -4x +5z =-17 A) x + z = 2 B) -4x +5z = -17 A) x = 2 -z then substituting this into B) B) -4* (2-z) + 5z = -17 -8 +4z + 5z = -17 9z = -9 z=-1 Substituting z = -1 into equation C 2x + 2z = 4 2x -2 = 4 x = 3